OpenStudy (lukecrayonz):

ln(x+2)+ln(x)=ln(35)

hartnn (hartnn):

you want to solve for 'x' , right ? use the property of log : \(\log A+\log B =\log AB\) on left side, so what you get for, ln(x+2)+ln(x) =...?

OpenStudy (lukecrayonz):

ln(2x+2)?

OpenStudy (lukecrayonz):

Or x^2+2x?

hartnn (hartnn):

x (x+2) = x*x + 2*x = x^2 +2x :) so you get ln (x^2+2x) = ln 35 so now can you solve the quadratic equation x^2+2x=35 ??

hartnn (hartnn):

ln(x+2)+ln(x) =ln 35 ---> x^2+2x=35 because if ln A =ln B, then A=B

hartnn (hartnn):

so ? could you calculate values of x ?